OpenStudy (anonymous):

2 out of 3 students in Mr. Pearson's ninth grade class recycle aluminum cans. If 48 recycle aluminum cans , how many students are in Mr.Pearson's class?

OpenStudy (kira_yamato):

3/2 x 48 = 72

OpenStudy (anonymous):

Omg that was my same question :0 Its 72 Students

OpenStudy (anonymous):

explain me that y we take 2/3X

OpenStudy (luffingsails):

The statement can be set up like this: (2/3) * (Number of Students in class) = 48 (=number of students who recycle) Now just solve for number of students in class (which is where multiplying by 3/2 comes in. (3/2)*(2/3)*(N) = 48*(3/2) N = (3/2)*48 N = 72 students